Find the length of a diagonal of a rectangle with l = 72 feet and w = 30 feet
Found 2 solutions by rfer, checkley77:
Answer by rfer(16322)

a^2+b^2=c^2
72^2+30^2=c^2
5184+900=6084
take sq rt of 6084
c=diagonal=78ft

Answer by checkley77(12844)

What you now have is a right triangle with sides of 72 & 30 ft.
Looking for the hypotenuse.
a^2+b^2=c^2
72^2+30^2=c^2
5,184+900=c62
c^2=6,084
c=sqrt6,084
c=78 ft. ans. for the diagonal(hypotenuse).